Find the value of C in the rolle's theorem, where f(x) =cos x and (-pie/2)<C<(pie/2).


1
Expert's answer
2021-03-31T14:07:20-0400

We have "f(x)=cosx" .

Since cosine function is continuous and differentiable everywhere.

Therefore "f(x)" is continuous on "[-\frac{\pi}{2},\frac{\pi}{2}]" and differentiable on "(-\frac{\pi}{2},\frac{\pi}{2})" .

Also , "f(-\frac{\pi}{2})=cos (-\frac{\pi}{2})=0" and "f(\frac{\pi}{2})=cos (\frac{\pi}{2})=0"

"\therefore f(-\frac{\pi}{2})=f(\frac{\pi}{2})" .

Thus all the conditions of Rolle's theorem are satisfied.

Then there exists at least one point "C\isin (-\frac{\pi}{2},\frac{\pi}{2})" for which "f'(C)=0" .

Now "f'(C)=0\implies-sinC=0"

"\implies C=0"

Hence the value of the "C" is .